OpenStudy (anonymous):

Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (johnweldon1993):

You mean \[\huge (\sqrt[3]{x^4})^5\] ?

OpenStudy (johnweldon1993):

If this is true... Remember that \[\huge (\sqrt[3]{x})^5 = x^{\frac{5}{3}}\] So if that is true \[\huge (\sqrt[3]{x^4})^5 = (x^4)^{\frac{5}{3}}\]